# Constraining the Luminosity Function and Delay-Time Distribution of Short Gamma-Ray Bursts for Multimessenger Gravitational-Wave Detection Rate Estimation

### v2

@@ Abstract
pubyear: 2025pagerange: Constraining the Luminosity Function and Delay-Time Distribution of Short Gamma-Ray Bursts for Multimessenger Gravitational-Wave Detection Rate EstimationConstraining the Luminosity Function and Delay-Time Distribution of Short Gamma-Ray Bursts for Multimessenger Gravitational-Wave Detection Rate Estimation
@@ 1 Introduction
With the derived LF and formation rate of sGRBs, we can further investigate the joint detection prospects of sGRBs and their GW counterparts using GRB and GW facilities. The success of joint GW and sGRB detections depends on the sensitivities and capabilities of both GW detectors and GRB missions. Currently, the fourth observation (O4) run of the International Gravitational Wave Network-4 (IGWN4), which includes four detectors—Advanced LIGO Livingston and Hanford (LIGO Scientific Collaboration et al., 2015), Advanced Virgo (Acernese et al., 2015), and KAGRA at design sensitivity (Aso et al., 2013)—is ongoing. In the future, the LIGO detectors at Livingston and Hanford will be upgraded to A$+$ sensitivity (Barsotti et al., 2018), and the planned GW detector LIGO-India (Iyer et al., 2011; Saleem et al., 2022) will joint the detector network, forming the International Gravitational Wave Network-5 (IGWN5). This IGWN5 network will significantly enhance the GW detection capabilities of its detectors. As the sensitivity and detection horizon of GW detectors continue to improve significantly, the primary limitation for joint detections is likely to soon be the sensitivity of GRB detectors such as Fermi. Recently, the Space-based multi-band astronomical Variable Objects Monitor (SVOM; Atteia et al. 2022) and the Einstein Probe (EP; Yuan et al. 2022) have been launched into space. Both satellites share the important scientific objective of detecting the electromagnetic counterparts of GWs. Estimating the expected detection rates of sGRB associated with GWs for these two missions will be crucial for subsequent observational plans and related research.
Several studies have aimed to estimate joint GW-sGRB detection rates. Howell et al. (2019) estimated that Advanced LIGO at design sensitivity could detect up to $\sim 4\,\mathrm{yr^{-1}}$ joint GW-sGRB events, increasing to $\sim 10\,\mathrm{yr^{-1}}$ with the A$+$ upgrade. For the fourth observing run (O4) of the LIGO-Virgo-KAGRA network, Colombo et al. (2022) projected joint detection rates of $\sim 0.03\,\mathrm{yr^{-1}}$ and $\sim 0.17\,\mathrm{yr^{-1}}$ using Swift and Fermi, respectively. Similarly, Patricelli et al. (2022) reported rates up to $\sim 6\,\mathrm{yr^{-1}}$ with Fermi during O4, while Salafia et al. (2023) predicted 0.2–1.3 detectable GW-sGRB events per year for Fermi/GBM and the Advanced GW detector network during O4. Bhattacharjee et al. (2024) found that for different jet structure models, joint GW-sGRB detection rates for Fermi and Swift with IGWN4 (IGWN5) range from 0.07–0.62 $\mathrm{yr^{-1}}$ (0.8–4.0 $\mathrm{yr^{-1}}$) and 0.02–0.14 $\mathrm{yr^{-1}}$ (0.15–1.0 $\mathrm{yr^{-1}}$), respectively.
@@
This paper is organized as follows. In Section 2, we describe the Fermi/GBM sample available to us, and obtain the pseudo redshifts of sGRBs. In Section 3, the LF and formation rate of sGRBs are determined through the maximum likelihood method. The synthetic BNS population model, the sGRB detection rates, and the prospects of joint sGRB and GW detections are presented in Section 4. Finally, we provide a brief summary and discussion in Section 5. Throughout this paper, we assume a standard $\Lambda$CDM cosmological model and adopt the parameters of $H_{0}=67.4$ $\mathrm{km}$ $\mathrm{s^{-1}}$ $\mathrm{Mpc^{-1}}$, $\mathrm{\Omega_{m}=0.315}$, and $\mathrm{\Omega_{\Lambda}=1-\mathrm{\Omega_{m}=0.315}}$ (Planck Collaboration et al., 2020).
@@ 2.2 Redshift estimation for sGRBs
The current number of sGRBs with redshift measurements is very small. In order to study their LF and formation rate, we determine the pseudo redshifts of sGRBs observed by Fermi/GBM using the empirical correlation between the observer-frame spectral peak energy ($E_{p}$) and the source-frame peak bolometric luminosity ($L$)(Tsutsui et al., 2013; Zhang & Wang, 2018). This correlation is given by (Zhang & Wang, 2018)
@@
where $d_{l}$ is the luminosity distance and $F_{p}$ is the source-frame peak bolometric flux at 64 ms time intervals(in units of $\mathrm{erg\,cm^{-2}\,s^{-1}}$). The pseudo redshift of each burst can be estimated using the spectral parameters of the Band function (Band et al., 1993) and the 64 ms peak photon flux $P$ in the 10–1000 keV observer-frame energy band. With the pseudo redshift, we can further calculate the peak bolometric luminosity $L$ in the 1–$10^{5}$ keV source-frame energy range. In our calculations, we adopt the spectral parameters from the Fermi/GBM burst catalog. For items that lack records, the low- and high-energy photon spectral indices are set to $\alpha=-1$ and $\beta=-2.25$ (Preece et al., 2000; Kaneko et al., 2006), and $E_{p}$ is randomly selected from the observed values. The pseudo redshift–luminosity distribution of 512 sGRBs is shown in Figure 2. Since the maximum redshift observed for sGRBs is 2.609, we exclude those bursts with pseudo-redshifts $z>3$. Hereafter, we use 478 sGRBs with $P\geq 4.1$ $\mathrm{photons\,cm^{-2}\,s^{-1}}$ and $z\leq 3$ for further analysis.
@@
Among our final sample of 478 sGRBs, 18 bursts have spectroscopically confirmed redshifts. In Table 1, we list the measured redshifts and pseudo-redshift estimates for these 18 bursts. To assess the consistency between the pseudo-redshift estimates and the observed redshifts, we perform a Kolmogorov-Smirnov (K-S) test. The K-S test statistic, $P_{\rm KS}$, quantifies the likelihood that two datasets are drawn from the same distribution: a higher $P_{\rm KS}$-value indicates greater consistency between the distributions. By convention, $P_{\rm KS}>0.05$ is considered insufficient evidence to reject the null hypothesis, while of $P_{\rm KS}<10^{-4}$ provides strong evidence against the null hypothesis (i.e., the distributions differ). The resulting $P_{\rm KS}$-value of 0.27 indicates that we cannot reject the null hypothesis. Therefore, we conclude that the redshift distribution estimated from the $E_{p}-L$ correlation is broadly consistent with the observed distribution, supporting the use of $E_{p}-L$ correlation as a statistically valid distance indicator for sGRBs.
@@ 3.1 Analysis method
where $f_{\rm sky}=0.7$ is the sky-coverage fraction of Fermi/GBM, [^2] Fermi is a low earth-orbit satellite that suffers from limited sky coverage because the Earth blocks $\sim 30\%$ of the sky. $f_{\rm not\_SAA}=0.85$ denotes the fraction of Fermi’s orbit outside the South Atlantic Anomaly (SAA), $R_{\rm sGRB}(z)$ is the comoving rate density of sGRBs in units of $\mathrm{Gpc^{-3}\,yr^{-1}}$, the factor $(1+z)^{-1}$ reflects the cosmological time dilation, $\phi(L)$ is the normalized LF of sGRBs, and $\mathrm{d}V(z)/\mathrm{d}z=4\pi cd_{l}^{2}(z)/[H_{0}(1+z)^{2}\sqrt{\Omega_{\mathrm{m}}(1+z)^{3}+\Omega_{\Lambda}}]$ is the comoving volume element in a flat $\Lambda$CDM cosmology.
GRB Name Measured Redshift Pseudo Redshift
GRB 090927422 1.37 0.66
GRB 100625773 0.45 0.66
GRB 200826187 0.75 0.25
GRB 090510016 0.90 1.14
GRB 080905499 0.12 1.82
GRB 231024556 0.06 1.31
GRB 150101641 0.13 0.26
GRB 131004904 0.72 0.79
GRB 100117879 0.92 1.01
GRB 100206563 0.41 0.55
GRB 111117510 2.21 0.69
GRB 210323918 0.42 0.87
GRB 160624477 0.48 1.55
GRB 100216422 0.04 0.28
GRB 220211047 0.53 0.36
GRB 201221963 1.05 0.53
GRB 210510806 0.22 1.11
GRB 160821937 0.16 0.59
Note.
It is believed that sGRBs originate from the mergers of two compact stars, either NS-NS or NS-BH. A compact binary system undergoes a long inspiral phase before the final merger, resulting in a significant time delay with respect to star formation (Faber & Rasio, 2012; Burns, 2020). If we assume that the fraction of mass forming compact binary systems relative to the total mass forming newborn stars remains constant over cosmological time, then the sGRB rate density $R_{\rm sGRB}(z)$ can be represented as a convolution of the star formation rate density (SFRD) $\psi_{\star}(z)$ and the probability density function of the delay time distribution $P(\tau)$ (Paul, 2018; Liu & Yu, 2019; Luo et al., 2022):
@@
where $\eta$ is the ‘$f_{b}$-free’ formation efficiency, still in units of $\mathrm{M^{-1}_{\odot}}$. Here, $z^{\prime}(\tau)$ denotes the redshift at the formation of the compact binary system, while $z$ corresponds to the redshift of the sGRB event. Thus, the delay time between the formation and merger of the compact binary system is given by $\tau=t(z)-t(z^{\prime})$, where $t(z)$ and $t(z^{\prime})$ are the ages of the Universe at redshifts $z$ and $z^{\prime}$, respectively. In view of the fact that the progenitors of NSs usually have initial masses of 10–29 $\mathrm{M_{\odot}}$, and their lifetimes are 10–30 $\mathrm{Myr}$, we set the minimum delay time to be $\tau_{\rm min}=10\,\mathrm{Myr}$, as done by Luo et al. (2022) in their treatment. The age of the Universe at redshift $z$, $t(z)$, is set as the maximum delay time $\tau_{\rm max}$. The SFRD $\psi_{\star}(z)$ can be expressed approximately by the analytical form (Yüksel et al., 2008):
@@ 3.2 Results
For a given delay-time distribution model, we optimize the free parameters by maximizing the likelihood function (Equation 3). The delay time distribution $P(\tau)$ has one or two free parameters (e.g., $t_{\rm G}$ and $\sigma_{\rm G}$ for Gaussian delay, $t_{\rm LN}$ and $\sigma_{\rm LN}$ for lognormal delay, or $\gamma$ for power-law delay). The sGRB LF $\phi(L)$ has three free parameters ($\nu_{1}$, $\nu_{2}$, and $L_{b}$). Additionally, the sGRB rate density $R_{\rm sGRB}(z)$, given by Equation (5), depends on the sGRB formation efficiency $\eta$, resulting in a total of five or six free parameters. For each model, the Python Markov chain Monte Carlo (MCMC) module, EMCEE (Foreman-Mackey et al., 2013), is employed to derive multidimensional parameter constraints from the observational data. The one-dimensional (1D) probability distributions and two-dimensional (2D) confidence regions (with $1\sigma-2\sigma$ contours) for the parameters in different delay-time distribution models are shown in Figure 3. These plots demonstrate that the sGRB formation efficiency ($\eta$) and the parameters of the sGRB LF ($\nu_{1}$, $\nu_{2}$, and $L_{b}$) are well constrained, regardless of the assumed delay-time distribution. For the power-law delay model, the power-law index ($\gamma$) is also well determined. In contrast, for the Gaussian and lognormal delay models, there is a strong degeneracy between the delay parameters: $t_{\rm G}$ and $\sigma_{\rm G}$ for the Gaussian model, and $t_{\rm LN}$ and $\sigma_{\rm LN}$ for the lognormal model. This degeneracy allows only an upper limit to be placed on $\sigma_{\rm G}$ (Gaussian) or $\sigma_{\rm LN}$ (lognormal).
Figure 3: Results of the MCMC inference for the observed sGRB sample. Panels (from left to right) show the 1D probability distributions and 2D confidence regions (with $1\sigma-2\sigma$ contours) for the parameters of the Gaussian, lognormal, and power-law delay-time distribution models.
@@
Gaussian $t_{\rm G}=1.14_{-0.17}^{+0.02}$,  $\sigma_{\rm G}<0.26$ $1.45_{-0.08}^{+0.12}$ $53.25_{-0.18}^{+0.13}$ $1.40_{-0.04}^{+0.04}$ $2.76_{-0.35}^{+0.64}$ $-47486$ 94996
Lognormal $t_{\rm LN}=2.14_{-0.28}^{+0.01}$,  $\sigma_{\rm LN}<0.06$ $1.30_{-0.06}^{+0.06}$ $53.22_{-0.10}^{+0.16}$ $1.37_{-0.03}^{+0.03}$ $2.72_{-0.28}^{+0.51}$ $-47520$ 95053
Power law $\gamma=2.04_{-0.48}^{+0.44}$ $1.51_{-0.06}^{+0.06}$ $53.17_{-0.09}^{+0.10}$ $1.29_{-0.04}^{+0.04}$ $2.63_{-0.30}^{+0.24}$ $-47526$ 95058
@@
With the best-fit parameters presented in Table 2, we are able to calculate the local sGRB formation rate at $z=0$, $R_{\mathrm{sGRB}}(0)$, using Equation (5). We find $R_{\mathrm{sGRB}}(0)=1.37_{-0.27}^{+0.30}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the Gaussian delay model, $1.63_{-0.27}^{+0.44}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the lognormal delay model, and $1.78_{-0.17}^{+0.12}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the power-law delay model, which are consistent with each other within the $1\sigma$ confidence level $1\sigma$. We would like to stress that our derived local sGRB formation rate is broadly consistent with previous results (Guetta & Piran, 2005; Yonetoku et al., 2014; Ghirlanda et al., 2016; Paul, 2018). The event rate of BNS mergers is estimated to be as high as $R_{\rm BNS}=320_{-240}^{+490}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$, as inferred from the second LIGO-Virgo GW transient catalog (Abbott et al., 2021). The discrepancy between the local sGRB formation rate and BNS merger rate arises primarily from the high collimation of GRB jets. For top-hat jets, the beaming factor is calculated as $f_{b}=R_{\mathrm{sGRB}}(0)/R_{\rm BNS}$, yielding a half-opening angle of $\theta_{j}\sim 3^{\circ}-12^{\circ}$. This range is in good agreement with previous estimates of $\theta_{j}\sim 3^{\circ}-26^{\circ}$ (Fong et al., 2013, 2015; Margutti et al., 2012).
[…]

### v3

@@ Abstract
pubyear: 2025pagerange: Constraining the Luminosity Function and Delay-Time Distribution of Short Gamma-Ray Bursts for Multimessenger Gravitational-Wave Detection Rate EstimationA
@@ 1 Introduction
With the derived LF and formation rate of sGRBs, we can further investigate the joint detection prospects of sGRBs and their GW counterparts using GRB and GW facilities. The success of joint GW and sGRB detections depends on the sensitivities and capabilities of both GW detectors and GRB missions. Currently, the fourth observation (O4) run of the International Gravitational Wave Network-4 (IGWN4), which includes four detectors—Advanced LIGO Livingston and Hanford (LIGO Scientific Collaboration et al., 2015), Advanced Virgo (Acernese et al., 2015), and KAGRA at design sensitivity (Aso et al., 2013)—is ongoing. In the future, the LIGO detectors at Livingston and Hanford will be upgraded to A++ sensitivity (Barsotti et al., 2018), and the planned GW detector LIGO-India (Iyer et al., 2011; Saleem et al., 2022) will joint the detector network, forming the International Gravitational Wave Network-5 (IGWN5). This IGWN5 network will significantly enhance the GW detection capabilities of its detectors. As the sensitivity and detection horizon of GW detectors continue to improve significantly, the primary limitation for joint detections is likely to soon be the sensitivity of GRB detectors such as Fermi. Recently, the Space-based multi-band astronomical Variable Objects Monitor (SVOM; Atteia et al. 2022) and the Einstein Probe (EP; Yuan et al. 2022) have been launched into space. Both satellites share the important scientific objective of detecting the electromagnetic counterparts of GWs. Estimating the expected detection rates of sGRB associated with GWs for these two missions will be crucial for subsequent observational plans and related research.
Several studies have aimed to estimate joint GW-sGRB detection rates. Howell et al. (2019) estimated that Advanced LIGO at design sensitivity could detect up to $\sim 4\,\mathrm{yr^{-1}}$ joint GW-sGRB events, increasing to $\sim 10\,\mathrm{yr^{-1}}$ with the A++ upgrade. For the fourth observing run (O4) of the LIGO-Virgo-KAGRA network, Colombo et al. (2022) projected joint detection rates of $\sim 0.03\,\mathrm{yr^{-1}}$ and $\sim 0.17\,\mathrm{yr^{-1}}$ using Swift and Fermi, respectively. Similarly, Patricelli et al. (2022) reported rates up to $\sim 6\,\mathrm{yr^{-1}}$ with Fermi during O4, while Salafia et al. (2023) predicted 0.2–1.3 detectable GW-sGRB events per year for Fermi/GBM and the Advanced GW detector network during O4. Bhattacharjee et al. (2024) found that for different jet structure models, joint GW-sGRB detection rates for Fermi and Swift with IGWN4 (IGWN5) range from 0.07–0.62 $\mathrm{yr^{-1}}$ (0.8–4.0 $\mathrm{yr^{-1}}$) and 0.02–0.14 $\mathrm{yr^{-1}}$ (0.15–1.0 $\mathrm{yr^{-1}}$), respectively.
@@
[…]
@@ 2.2 Redshift estimation for sGRBs
The current number of sGRBs with redshift measurements is very small. In order to study their LF and formation rate, we determine the pseudo redshifts of sGRBs observed by Fermi/GBM using the empirical correlation between the observer-frame spectral peak energy ($E_{p}$) and the source-frame peak bolometric luminosity (LL)(Tsutsui et al., 2013; Zhang & Wang, 2018). This correlation is given by (Zhang & Wang, 2018)
@@
[…]
@@
Figure 3: Normalized cumulative redshift distribution for 18 sGRBs with known redshifts in our sample (blue line). The orange line represents the pseudo-redshift distribution for these bursts derived from the $E_{p}-L$ correlation. The gray lines show uncertainty regions estimated from 100 Monte Carlo simulations.
Among our final sample of 478 sGRBs, 18 bursts have spectroscopically confirmed redshifts. In Table 1, we list the measured redshifts and pseudo-redshift estimates for these 18 bursts. Note that among these 18 bursts, only GRB 210510806 lacks a measured $E_{p}$. Following Paul (2018a, b), we randomly sample this burst’s $E_{p}$ from the observed $E_{p}$ distribution of Fermi sGRBs with spectral measurements. To validate our pseudo-redshift estimates, we compare the cumulative distributions of measured (blue line) and pseudo-redshifts (orange line) in Figure 3 and perform a Kolmogorov-Smirnov (K-S) test between these distributions. The K-S test statistic, $P_{\rm KS}$, quantifies the likelihood that two datasets are drawn from the same distribution: a higher $P_{\rm KS}$-value indicates greater consistency between the distributions. By convention, $P_{\rm KS}>0.05$ is considered insufficient evidence to reject the null hypothesis, while of $P_{\rm KS}<10^{-4}$ provides strong evidence against the null hypothesis (i.e., the distributions differ). The resulting $P_{\rm KS}$-value of 0.80 indicates that we cannot reject the null hypothesis. Moreover, we estimate an uncertainty region for the pseudo-redshift distribution of these 18 bursts. Redshift errors primarily originate from $E_{p}$ uncertainties, so we performed 100 Monte Carlo simulations per data point to derive cumulative distributions. Gray lines in Figure 3 show these results, indicating the uncertainty region partially covers the observed distribution (blue line). Therefore, we conclude that the redshift distribution estimated from the $E_{p}-L$ correlation is broadly consistent with the observed distribution, supporting the use of $E_{p}-L$ correlation as a statistically valid distance indicator for sGRBs. Notably, Paul (2018a) demonstrated that random $E_{p}$ sampling statistically replicates the redshift distribution of long GRBs. Figure 3 confirms this approach also works for sGRBs: their redshift distribution is replicated through $E_{p}$ uncertainty propagation and random $E_{p}$ draws for spectrally unmeasured bursts.
In our pseudo-redshift estimation, we make no assertion regarding the accuracy of individual values. We do not assess pseudo-redshift accuracy per GRB, as our analysis relies exclusively on statistical distribution comparisons. This approach assigns pseudo redshifts to bursts in a collective statistical context only. Nevertheless, this limitation does not invalidate our methodology, since we use pseudo-redshifts and derived luminosities solely as ensemble quantities for modeling the sGRB LF and formation rate.
GRB Name Measured Redshifta Pseudo Redshiftb
GRB 090927422 1.37 0.66
GRB 100625773 0.45 0.66
GRB 200826187 0.75 0.25
GRB 090510016 0.90 1.14
GRB 080905499 0.12 1.82
GRB 231024556 0.06 0.40
GRB 150101641 0.13 0.26
GRB 131004904 0.72 0.79
GRB 100117879 0.92 1.01
GRB 100206563 0.41 0.55
GRB 111117510 2.21 0.69
GRB 210323918 0.42 0.91
GRB 160624477 0.48 1.55
GRB 100216422 0.04 0.28
GRB 220211047 0.53 0.36
GRB 201221963 1.05 0.91
GRB 210510806 0.22 1.11c
GRB 160821937 0.16 0.59
Note.
aThe measured redshifts are obtained from https://user-web.icecube.wisc.edu/˜grbweb_public/Summary_table.html. bPseudo-redshift estimation involves statistical analysis, and measured versus pseudo redshifts are not meant to be equal since pseudo redshifts derive from the empirical $E_{p}$-$L$ correlation and, for bursts without spectral measurements, random $E_{p}$ sampling. cFor GRB 210510806, its reported pseudo redshift corresponds to a single random $E_{p}$ draw.
@@ 3.1 Analysis method
[…]
It is believed that sGRBs originate from the mergers of two compact stars, either NS-NS or NS-BH. A compact binary system undergoes a long inspiral phase before the final merger, resulting in a significant time delay with respect to star formation (Faber & Rasio, 2012; Burns, 2020). If we assume that the fraction of mass forming compact binary systems relative to the total mass forming newborn stars remains constant over cosmological time, then the sGRB rate density $R_{\rm sGRB}(z)$ can be represented as a convolution of the star formation rate density (SFRD) $\psi_{\star}(z)$ and the probability density function of the delay time distribution $P(\tau)$ (Paul, 2018b; Liu & Yu, 2019; Luo et al., 2022):
@@
where $\eta$ is the $f_{b}$-free’ formation efficiency, still in units of $\mathrm{M^{-1}_{\odot}}$. Here, $z^{\prime}(\tau)$ denotes the redshift at the formation of the compact binary system, while $z$ corresponds to the redshift of the sGRB event. Thus, the delay time between the formation and merger of the compact binary system is given by $\tau=t(z)-t(z^{\prime})$, where $t(z)$ and $t(z^{\prime})$ are the ages of the Universe at redshifts $z$ and $z^{\prime}$, respectively. In view of the fact that the progenitors of NSs usually have initial masses of 10–29 $\mathrm{M_{\odot}}$, and their lifetimes are 10–30 $\mathrm{Myr}$, we set the minimum delay time to be $\tau_{\rm min}=10\,\mathrm{Myr}$, as done by Luo et al. (2022) in their treatment. The age of the Universe at redshift $z$, $t(z)$, is set as the maximum delay time $\tau_{\rm max}$. The SFRD $\psi_{\star}(z)$ can be expressed approximately by the analytical form (Yüksel et al., 2008):
@@ 3.2 Results
[…]
@@
Gaussian $t_{\rm G}=1.14_{-0.17}^{+0.02}$,  $\sigma_{\rm G}<0.26$ $1.45_{-0.08}^{+0.12}$ $53.25_{-0.18}^{+0.13}$ $1.40_{-0.04}^{+0.04}$ $2.76_{-0.35}^{+0.64}$ 47486-47486 94996
Lognormal $t_{\rm LN}=2.14_{-0.28}^{+0.01}$,  $\sigma_{\rm LN}<0.06$ $1.30_{-0.06}^{+0.06}$ $53.22_{-0.10}^{+0.16}$ $1.37_{-0.03}^{+0.03}$ $2.72_{-0.28}^{+0.51}$ 47520-47520 95053
Power law $\gamma=2.04_{-0.48}^{+0.44}$ $1.51_{-0.06}^{+0.06}$ $53.17_{-0.09}^{+0.10}$ $1.29_{-0.04}^{+0.04}$ $2.63_{-0.30}^{+0.24}$ 47526-47526 95058
@@
With the best-fit parameters presented in Table 2, we are able to calculate the local sGRB formation rate at $z=0$, $R_{\mathrm{sGRB}}(0)$, using Equation (5). We find $R_{\mathrm{sGRB}}(0)=1.37_{-0.27}^{+0.30}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the Gaussian delay model, $1.63_{-0.27}^{+0.44}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the lognormal delay model, and $1.78_{-0.17}^{+0.12}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the power-law delay model, which are consistent with each other within the $1\sigma$ confidence level $1\sigma$. We would like to stress that our derived local sGRB formation rate is broadly consistent with previous results (Guetta & Piran, 2005; Yonetoku et al., 2014; Ghirlanda et al., 2016; Paul, 2018b). The event rate of BNS mergers is estimated to be as high as $R_{\rm BNS}=320_{-240}^{+490}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$, as inferred from the second LIGO-Virgo GW transient catalog (Abbott et al., 2021). The discrepancy between the local sGRB formation rate and BNS merger rate arises primarily from the high collimation of GRB jets. For top-hat jets, the beaming factor is calculated as $f_{b}=R_{\mathrm{sGRB}}(0)/R_{\rm BNS}$, yielding a half-opening angle of $\theta_{j}\sim 3^{\circ}-12^{\circ}$. This range is in good agreement with previous estimates of $\theta_{j}\sim 3^{\circ}-26^{\circ}$ (Fong et al., 2013, 2015; Margutti et al., 2012).